OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (nurali):

−4|x + 5| = −16 Divide by -4 |x + 5| = 4 x+5=4 or x+5=-4 x=4-5 or x=-4-5 x=-1 or x=-9

OpenStudy (anonymous):

whats the answer?

OpenStudy (nurali):

Third option